# TS-RAG: Retrieval-Augmented Generation based Time Series Foundation Models are Stronger Zero-Shot Forecaster

### v2

@@ Abstract
Recently, Large Language Models (LLMs) and Foundation Models (FMs) have become prevalent for time series forecasting tasks. However, fine-tuning large language models (LLMs) for forecasting enables the adaptation to specific domains but may not generalize well across diverse, unseen datasets. Meanwhile, existing time series foundation models (TSFMs) lack inherent mechanisms for domain adaptation and suffer from limited interpretability, making them suboptimal for zero-shot forecasting. To this end, we present TS-RAG, a retrieval-augmented generation based time series forecasting framework that enhances the generalization capability and interpretability of TSFMs. Specifically, TS-RAG leverages pre-trained time series encoders to retrieve semantically relevant time series segments from a dedicated knowledge database, incorporating contextual patterns for the given time series query. Next, we develop a learnable Mixture-of-Experts (MoE)-based augmentation module, which dynamically fuses retrieved time series patterns with the TSFM’s representation of the input query, improving forecasting accuracy without requiring task-specific fine-tuning. Thorough empirical studies on seven public benchmark datasets demonstrate that TS-RAG achieves state-of-the-art zero-shot forecasting performance, outperforming TSFMs by up to 6.51% across diverse domains and showcasing desired interpretability.
Time Series Forecasting, Time Series Foundation Model, RAG, Zero-shot
@@ 1 Introduction
Time series forecasting, which aims to predict future values of a sequence based on its past observations, plays a critical role in various real-world applications, e.g., finance, healthcare, energy management, and climate science. The key idea is to capture the temporal dependency patterns in the form of trend, seasonality, autocorrelation, etc, to make accurate predictions and generalize across different datasets.
@@
In the past, a substantial amount of effort has been made to tackle this problem. Traditional statistical methods such as AutoRegressive Integrated Moving Average (ARIMA) (Whittle, 1951) work well for stationary time series but struggle with complex dependencies and non-linear patterns. Machine learning approaches, such as Random Forest (Breiman, 2001) and XGBoost (Chen & Guestrin, 2016), can handle external covariates of features but fail to capture long-range dependencies. Deep learning techniques, including Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997), Gated Recurrent Units (GRUs) (Cho et al., 2014), Temporal Convolutional Networks (TCNs) (Lea et al., 2016), Graph Neural Networks (GNN) based models (Cao et al., 2020; Shang et al., 2021), and transformer based models (Vaswani et al., 2017; Zhou et al., 2021) are typically trained on specific domains and may not perform well to diverse unseen datasets. More recently, there is a prevalent interest in adapting Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023b; Achiam & et al., 2023) for time series tasks (Gruver et al., 2024a; Jiang et al., 2024) and developing Foundation Models (FM) tailored for time series data (Rasul et al., 2023; Garza & Mergenthaler-Canseco, 2023; Liu et al., 2024b). Although both LLMs and FM have shown great promise in improving forecasting accuracy and handling complex temporal dynamics, they still face immense barriers when applied to zero-shot forecasting tasks, limiting their real-world applicability.
@@
Specifically, fine-tuning LLMs for time series forecasting can enable the adaptation to specific datasets, however, those methods could struggle with generalization across diverse, unseen domains (Zhou et al., 2023; Jin et al., 2023; Pan et al., 2024). In addition, they typically involve heavy computational costs, even with limited samples. Recent work (Gruver et al., 2024b) has explored leveraging LLMs as zero-shot forecasters. Time Series Foundation Models (TSFMs)  (Das et al., 2023; Ekambaram et al., 2024; Ansari et al., 2024; Woo et al., 2024), while effective in learning general time series representations, lack inherent mechanisms for domain adaptation as they cannot incorporate external contextual knowledge dynamically, making them less robust when faced with complex and evolving time series patterns. Furthermore, TSFMs often suffer from limited interpretability.
@@
Recently, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) has demonstrated significant success across various Natural Language Processing (NLP) tasks by enhancing LLMs through the retrieval of relevant document segments from external knowledge bases. By incorporating retrieved information, RAG refines the existing prompts and generates more informed, context-aware outputs, improving both accuracy and adaptability in diverse applications. Inspired by this, in this paper, we present TS-RAG, a retrieval-augmented generation based time series forecasting framework to dynamically incorporate semantically relevant time series patterns into its forecasting pipeline, eliminating the need for fine-tuning while significantly improving zero-shot forecasting performance and the interpretability of TSFMs.
@@
Instead of simply relying on the input time series query, TS-RAG first adopts pre-trained time series encoders to retrieve relevant time series segments from a dedicated knowledge database, providing valuable contextual knowledge for forecasting. Next, to effectively integrate retrieved time series knowledge, TS-RAG leverages a learnable Mixture-of-Experts (MoE)-based (Shazeer et al., 2017) augmentation module which can dynamically fuse retrieved patterns with the input time series query, ensuring that the model benefits from both existing knowledge and current query. With retrieval-augmented generation, TS-RAG not only can circumvent the need for fine-tuning on specific datasets but also can utilize retrieved segments to provide explicit rationales to explain the model’s predictions. Finally, thorough empirical studies on seven public benchmark datasets demonstrate that TS-RAG achieves state-of-the-art zero-shot forecasting performance, outperforming existing TSFMs by up to 6.51% across diverse domains while simultaneously enhancing interpretability, reinforcing its potential as a robust and generalizable forecasting framework.
@@ 2 Related Work
2.1 Time Series Foundation Models
Recently, the rapid development of Time Series Foundation Models has drawn significant attention and made substantial progress in time series forecasting. These existing models, often adapted from advancements in Natural Language Processing (NLP) and Vision Transformers, have demonstrated strong generalization capabilities across diverse datasets. Lag-Llama (Rasul et al., 2023) and TimeGPT-1 (Garza & Mergenthaler-Canseco, 2023) are pioneering forecasting foundation models, pre-trained on extensive time series datasets spanning multiple domains. Lag-Llama utilizes lagged time series features and the LLaMA architecture (Touvron et al., 2023a), while TimeGPT-1 adopts an encoder-decoder transformer structure to handle forecasting tasks effectively. Tiny Time Mixers (TTMs) (Ekambaram et al., 2024), built upon the earlier work TSMixer, trains a compact foundation model using multi-resolution data from various domains. TimesFM (Das et al., 2023) pretrains a patched-decoder attention model on a large time-series corpus to enable zero-shot forecasting across multiple domains. Chronos  (Ansari et al., 2024) is an encoder-decoder style probabilistic time series foundation model that employs next-token-prediction for forecasting, enabling it to perform zero-shot forecasting on unseen forecast tasks. As a subsequent version of Chronos, Chronos-bolt incorporates a patch-based input strategy and uses decoder representations to generate quantile forecasts across multiple future steps, further improving forecast accuracy over its predecessor. Moirai (Woo et al., 2024) introduces a masked encoder-based universal time series forecasting transformer, accompanied by a novel large-scale time series dataset, LOSTA. Similarly, MOMENT (Goswami et al., 2024) compiles a large and diverse collection of public time series, called the Time series Pile, and systematically tackles the large-scale multi-dataset pretraining problem. These methods, however, lack inherent mechanisms to incorporate external contextual knowledge dynamically to facilitate zero-shot learning and suffer from limited interpretability.
2.2 Retrieval-Augmented for Time Series Forecasting
Existing LLM-based time series forecasters (Zhou et al., 2023; Jin et al., 2023; Pan et al., 2024) have demonstrated remarkable achievements in in-domain time series analysis. Nonetheless, adapting these models to different domains requires substantial computational resources. Given the challenges of adapting LLM-based models across different domains due to the computational costs, not only do the previously mentioned time series foundation models offer a lightweight solution to this issue but also the advent of Retrieval-Augmented Generation (RAG) techniques  (Lewis et al., 2020) presents a compelling alternative. By augmenting the generative model’s input with retrieved external knowledge, RAG has proven effective in open-domain question answering and document generation in the language field. Applying RAG to time series forecasting represents a novel and pioneering research direction. ReTime (Jing et al., 2022) proposes relational retrieval and content synthesis for spatial-temporal time series and time series imputation analysis.
TimeRAG (Yang et al., 2024) leverages K-means clustering to build a time series knowledge base and employs Dynamic Time Warping (DTW) as a similarity metric for retrieval. RATD (Liu et al., 2024a) utilizes retrieved historical time series to guide the denoising process of the diffusion model. RAF (Tire et al., 2024) proposes concatenating the retrieved context with the original input to create an augmented input for forecasting. However, existing works fail to unleash the full potential of RAG in time series analysis. Our proposed TS-RAG is specially designed to enhance zero-shot forecasting in TSFMs, incorporating an innovative learnable Mixture-of-Experts (MoE) augmentation module that can dynamically fuse retrieved patterns with the input time series query.
@@ 3 TS-RAG for Zero-Shot Time Series Forecasting
Figure 1: The TS-RAG model architecture processes an input time series by retrieving the top-$k$ similar time series and corresponding future horizons from a knowledge base through embedding similarity. These elements are fused with the input series embeddings using our MOE architecture to generate the final time series forecast.
@@
Overview: The model architecture of the proposed TS-RAG consists of three key components, i.e., TSFM, retriever, and augmentation module, as shown in Figure 1. Given an input time series context window, a pretrained retrieval encoder first generates the context embedding. This embedding is then compared with time series context embeddings previously stored in the retrieval knowledge base to retrieve the top-$k$ similar time series pairs. Each retrieved pair includes a historical context and its corresponding forecasting horizon, which are utilized for augmentation and fusion to refine the zero-shot time series forecasting.
@@
The retrieved future horizons of top–$k$ similar time series pairs are first transformed into embeddings and then fed into the Mixture-of-Experts (MoE) augmentation module along with the input time series embedding generated by the TSFM backbone. The MoE module adaptively assigns importance scores to these embeddings, dynamically integrating them into a unified representation. This final representation is then passed through the output projection layer of the TSFM to produce the enhanced time series forecast.
We provide detailed descriptions of the retrieval knowledge base, the TS-RAG framework, and the adaptive pretraining with zero-shot inference in the following sections.
3.1 Retrieval Knowledge Base for TS-RAG
TSFMs are typically pretrained on a multi-domain time series dataset, enabling them to perform zero-shot forecasting in unseen scenarios. Similarly, to enhance the generalization capability of TS-RAG, we construct a multi-domain dataset for its adaptive pretraining, specifically designed to learn the MoE module.
@@ 3.1 Retrieval Knowledge Base for TS-RAG
We leverage the pretraining dataset of Chronos (Ansari et al., 2024), which utilizes TSMixup to randomly combine time series data points from various domains. This approach enhances data diversity by blending different patterns, thereby improving the model’s ability to generalize. Given that the additional parameters of TS-RAG are significantly less than those in the TSFM backbone, we sample a much smaller subset from the Chronos pretraining dataset to serve as the adaptive pretraining dataset for TS-RAG. From this smaller subset, we selectively extract a further subset to construct the retrieval knowledge database for TS-RAG, which will be used in the forecasting phase.
@@
The time series data stored in the knowledge base is processed into standard pairs, each consisting of a context window and its corresponding forecasting horizon. Formally, this can be expressed as: $\{(x_{i},y_{i})\,|\,i=1,2,\ldots,n\}$ where $x_{i}$ is the context window of the $i$-th time series, $y_{i}$ is the future horizon associated with $x_{i}$ and $n$ is the total number of pairs in the knowledge base.
Additionally, we employ a pretrained retrieval encoder to generate embeddings for the time series contexts stored in the retrieval knowledge base. These embeddings are then stored along with the corresponding time series data within the knowledge base. As a result, the structure of the retrieval knowledge base can be formally described as the following set of triplets:
$\mathcal{D}=\{(x_{i},e_{i},y_{i})\,|\,i=1,2,\ldots,n\}$
@@
where $\mathcal{D}$ denotes the retrieval knowledge base and $e_{i}$ denotes the embedding of $x_{i}$ generated by the pretrained retrieval encoder, resulting in a highly efficient training process.
3.2 Retrieve-Augmented Generation based Time Series Foundation Models

### v3

@@ Abstract
Large Language Models (LLMs) and Foundation Models (FMs) have recently become prevalent for time series forecasting tasks. While fine-tuning LLMs enables domain adaptation, they often struggle to generalize across diverse and unseen datasets. Moreover, existing Time Series Foundation Models (TSFMs) still face challenges in handling non-stationary dynamics and distribution shifts, largely due to the lack of effective mechanisms for adaptation. To this end, we present TS-RAG, a retrieval-augmented generation framework for time series forecasting that enhances the generalization and interpretability of TSFMs. Specifically, TS-RAG leverages pre-trained time series encoders to retrieve semantically relevant segments from a dedicated knowledge base, enriching the contextual representation of the input query. Furthermore, we propose an Adaptive Retrieval Mixer (ARM) module that dynamically fuses the retrieved patterns with the TSFM’s internal representation, improving forecasting accuracy without requiring task-specific fine-tuning. Thorough empirical studies on seven public benchmark datasets demonstrate that TS-RAG achieves state-of-the-art zero-shot forecasting performance, outperforming the existing TSFMs by up to 6.84% across diverse domains while also providing desirable interpretability.
footnotetext: Correspondence to: Lintao Ma <lintao.mlt@antgroup.com>, Yuriy Nevmyvaka <yuriy.nevmyvaka@- morganstanley.com>, Dongjin Song <dongjin.song@uconn.edu>.
@@ 1 Introduction
Time series forecasting, which aims to predict future values of a sequence based on its past observations, plays a critical role in various real-world applications, e.g., finance, healthcare, energy management, and climate science. Modeling time series data essentially captures the temporal dependency patterns in the form of trend, seasonality, autocorrelation, etc, to make accurate predictions and generalize across different datasets. In the past, a substantial amount of effort has been made to tackle this problem. Traditional statistical methods such as AutoRegressive Integrated Moving Average (ARIMA) [1] work well for stationary time series but struggle with complex dependencies and non-linear patterns. Machine learning approaches, such as Random Forest [2] and XGBoost [3], can handle external covariates of features but fail to capture long-range dependencies. Deep learning techniques, including Long Short-Term Memory (LSTM) [4], Gated Recurrent Units (GRUs) [5], Temporal Convolutional Networks (TCNs) [6], Graph Neural Networks (GNN) based models [7, 8], and transformer based models [9, 10] are typically trained within specific domains and often struggle to generalize effectively to diverse, unseen datasets. More recently, there is a prevalent interest in adapting Large Language Models (LLMs) [11, 12, 13] for time series tasks [14, 15] and developing Foundation Models (FM) tailored for time series data [16, 17, 18]. Although both LLMs and FM have shown great promise in improving forecasting accuracy and handling complex temporal dynamics, they still face immense barriers when applied to zero-shot forecasting tasks, limiting their real-world applicability.
@@
Specifically, recent work [19] has explored the potential usage of LLMs as zero-shot forecasters for time series tasks. While fine-tuning LLMs can facilitate adaptation and yield better performance over specific datasets, such methods often struggle to generalize across diverse, unseen domains [20, 21, 22], and typically involve substantial computational overhead, even when applied to limited data. Time Series Foundation Models (TSFMs)  [23, 24, 25, 26] have emerged as a promising alternative by learning time series representations across a wide range of time series data. However, they lack inherent mechanisms for domain adaptation, as they cannot incorporate external contextual knowledge dynamically. This shortcoming limits their capability in handling complex, non-stationary, and evolving time series patterns. Furthermore, most TSFMs offer limited interpretability, which poses challenges for their deployment in high-stakes or decision-critical applications.
@@
Recently, Retrieval-Augmented Generation (RAG) [27] has demonstrated significant success across various Natural Language Processing (NLP) tasks [28, 29, 30, 31]. By retrieving relevant document segments and incorporating them into prompts, RAG refines the existing prompts and enables LLMs to generate more informed, context-aware outputs, thereby improving both accuracy and adaptability in diverse applications. Motivated by this paradigm, we propose TS-RAG, a retrieval-augmented generation based time series forecasting framework. TS-RAG dynamically retrieves semantically relevant time series patterns and integrates them into the forecasting pipeline, enabling strong zero-shot performance without the need for fine-tuning. In addition, it significantly enhances the interpretability of TSFMs by providing contextual evidence for their predictions.
@@
Instead of simply relying on the input time series query, TS-RAG first adopts pre-trained time series encoders to retrieve relevant time series segments from a dedicated knowledge database, providing valuable contextual knowledge for forecasting. Next, to effectively integrate retrieved time series knowledge, TS-RAG leverages a learnable Adaptive Retrieval Mixer (ARM) augmentation module, which can dynamically fuse retrieved patterns with the input time series query, ensuring that the model benefits from both existing knowledge and the current query. With retrieval-augmented generation, TS-RAG not only can circumvent the need for fine-tuning on specific datasets but also can utilize retrieved segments to provide explicit rationales to enhance the interpretability of the model’s predictions. Finally, thorough empirical studies on seven public benchmark datasets demonstrate that TS-RAG achieves state-of-the-art zero-shot forecasting performance, outperforming existing TSFMs by up to 6.84% across diverse domains while exhibiting desirable interpretability, highlighting its potential as a robust and generalizable forecasting framework.
@@ 2 Related Work
Time Series Foundation Models Existing LLM-based time series forecasters [20, 32, 33, 21, 34, 22] have demonstrated remarkable achievements in in-domain time series analysis. However, the domain adaptation challenges and significant computational costs associated with LLM-based models have motivated the emergence of time series foundation models as a more efficient and scalable alternative. Inspired by recent advancements in Natural Language Processing (NLP) and Vision Transformers [35], Time Series Foundation Models have rapidly developed and drawn significant attention. These models have demonstrated strong generalization capabilities across diverse datasets, leading to substantial progress in time series forecasting. Lag-Llama [16] and TimeGPT-1 [17] are pioneering forecasting foundation models, pre-trained on extensive time series datasets spanning multiple domains. Lag-Llama utilizes lagged time series features and the LLaMA architecture [36], while TimeGPT-1 adopts an encoder-decoder transformer structure to handle forecasting tasks effectively. Following the paradigm of patching tokenization and optimization, models such as TimesFM [23], MOMENT [37], Timer [38, 39], and Sundial [40] first patch and embed continuous time series values, and subsequently model the output distributions and perform point forecasting. To further improve efficiency, Tiny Time Mixers (TTMs)[24] train a compact foundation model, while TimeMOE[41] adopts a sparse mixture-of-experts architecture that activates only a subset of expert networks during inference, thereby maintaining strong performance with reduced computational cost. However, deterministic predictions usually cannot satisfy the requirement of decision-making. To address this limitation, Moirai [26] trains a probabilistic model that captures a mixture of distributions. Building on language modeling techniques, Chronos [25] discretizes time series through scaling and quantization, and models the resulting categorical distributions using cross-entropy loss. To enhance inference efficiency and forecasting accuracy, Chronos-bolt replaces discrete tokenization with a patch-based input strategy and utilizes decoder representations to produce quantile forecasts over multiple future steps, achieving improved performance over its predecessor. These methods, however, lack inherent mechanisms to incorporate external contextual knowledge dynamically to facilitate zero-shot learning and suffer from limited interpretability.
Retrieval-Augmented for Time Series Forecasting Although both LLMs and TSFMs have achieved strong performance in time series forecasting, they can still struggle in scenarios involving non-stationarity or distribution shifts. To address these challenges, Retrieval-Augmented Generation (RAG) techniques [27] offer a promising approach to incorporate external knowledge and enhance generalization and robustness. Recent works have explored the integration of retrieval mechanisms into time series forecasting. Among them, several approaches rely on fine-tuning to adapt the model to downstream tasks, including ReTime [42], RATD [43], TimeRAG [44], and RAFT [45]. ReTime [42] proposes relational retrieval and content synthesis for retrieval-based time series forecasting. RATD [43] utilizes retrieved historical time series to guide the denoising process of diffusion models, enhancing forecasting performance. RAFT [45] integrates retrieval with a multi-resolution forecasting framework. These methods typically construct the retrieval database from the training set of the target task and require fine-tuning for model adaptation. Similarly, TimeRAG [44] incorporates retrieved sequences into the forecasting process by using a frozen LLM backbone, and introduces a trainable reprogramming layer to align time series and text modalities.
RAF [46] is a pioneering work that introduces a retrieval-augmented framework for zero-shot time series forecasting. It utilizes Chronos as the backbone model and constructs the augmented input by directly concatenating the processed retrieved context with the original time series input. However, this approach may face scalability and efficiency challenges when applied to large-scale retrieval databases. Moreover, RAG techniques for time series forecasting remain underexplored, particularly in terms of knowledge base construction, augmentation strategies, and their integration with TSFMs to facilitate zero-shot forecasting. To bridge these gaps, we propose TS-RAG, a retrieval-augmented framework specifically designed to enhance zero-shot forecasting capabilities of TSFMs.
@@ 3 TS-RAG for Zero-Shot Time Series Forecasting
Overview: The proposed TS-RAG consists of three key components, i.e., a time series foundation model, a retriever, and a learnable Adaptive Retrieval Mixer (ARM) augmentation module, as shown in Figure 1. Given an input time series, a pretrained retriever encoder first generates the corresponding embedding. This embedding is then compared with time series context embeddings previously stored in the retrieval knowledge base to retrieve the top-$k$ similar time series pairs. Each retrieved pair includes a historical context and its corresponding forecasting horizon, and the forecasting horizon is utilized for augmentation to refine the zero-shot time series forecasting.
Figure 1: The TS-RAG model architecture processes an input time series by retrieving the top-$k$ semantically similar time series segments and their corresponding future horizons from a knowledge base (via Retriever), based on embedding similarity. These retrieved segments are then integrated with the input series embedding using the proposed Adaptive Retrieval Mixer (ARM) augmentation module, enabling the model to generate the final forecast with enriched contextual information.
@@
The retrieved future horizons of top-$k$ similar time series pairs are first transformed into embeddings and then fed into the ARM augmentation module along with the input time series embedding generated by the TSFM backbone. The ARM augmentation module adaptively assigns importance scores to these embeddings, dynamically integrating them into a unified representation. This final representation is then passed through the output projection layer of the TSFM to produce the enhanced time series forecast. Below, we provide details for the construction of the retrieval knowledge base (Section 3.1), the TS-RAG framework architecture (Section 3.2), and the pretraining strategy used to enable zero-shot inference (Section 3.3).
3.1 Construction of the Retrieval Knowledge Base
@@
TSFMs are typically pretrained on a multi-domain time series dataset to enhance their generalization capability. For TS-RAG, we adopt a similar strategy, i.e., construct a multi-domain dataset and specifically focus on learning the ARM augmentation module. We leverage the pretraining dataset of Chronos [25], which utilizes TSMixup to randomly combine time series data points from various domains. This approach enhances data diversity by blending different patterns, thereby improving the model’s ability to generalize. Given the fact that the trainable parameters of TS-RAG are significantly less than those in the TSFM backbone and the retrieval encoder (more discussions are available in Appendix B.5), we uniformly sample a subset from the Chronos pretraining dataset to serve as the pretraining dataset for TS-RAG. Based on this subset, we can further draw a subset to construct the retrieval knowledge database for TS-RAG, which will be used in the inference (forecasting) stage.
@@
The time series data stored in the knowledge base is processed into standard pairs, each consisting of a context window and its corresponding forecasting horizon. Formally, this can be expressed as: $\{(\mathbf{x}_{i},\mathbf{y}_{i})\,|\,i=1,2,\ldots,n\}$ where $\mathbf{x}_{i}\in\mathbb{R}^{T}$ is the context window of the $i$-th time series with length $T$, $\mathbf{y}_{i}\in\mathbb{R}^{L}$ is the future horizon of length $L$ associated with $\mathbf{x}_{i}$, and $n$ is the total number of pairs in the knowledge base. Based on the time series context windows $\{\mathbf{x}_{i}|i=1,2,\ldots n\}$ stored in the retrieval knowledge base, we employ a pretrained retrieval encoder to generate their embeddings $\{\mathbf{e}_{i}|\,i=1,2,\ldots n\}$ where $\mathbf{e}_{i}\in\mathbb{R}^{d}$. These embeddings are then stored along with the corresponding time series data in the knowledge base. As a result, the structure of the retrieval knowledge base can be formally described as the following set of triplets:
$\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{e}_{i},\mathbf{y}_{i}),|\,i=1,2,\ldots,n\}$ (1)
@@
where $\mathcal{D}$ represents the retrieval knowledge base, and $\mathbf{e}_{i}$ denotes the embedding of the input sequence $\mathbf{x}_{i}$, computed by a pretrained retrieval encoder, thereby facilitating an efficient training process.
3.2 Architecture of TS-RAG Framework